OpenStudy (sissyedgar):

Question Below!!! @misty1212

OpenStudy (sissyedgar):

:)

OpenStudy (misty1212):

lol i totally answered this before!

OpenStudy (sissyedgar):

you did? o, i lost it! :P

OpenStudy (misty1212):

one mile is 63360 inches so one inch is \(\frac{1}{63360}\) miles lets put in in a calculator and then write it in scientific notation

OpenStudy (misty1212):

http://www.wolframalpha.com/input/?i=1%2F63360

OpenStudy (sissyedgar):

Its not in scientific notation tho

OpenStudy (misty1212):

\[0.00001578\] in scientific notation \[\huge 1.578\times 10^{-5}\]

OpenStudy (sissyedgar):

o, lml

OpenStudy (misty1212):

i moved the decimal and counted the spaces since it went left, the exponent is negative

OpenStudy (misty1212):

i mean right

OpenStudy (sissyedgar):

right

OpenStudy (misty1212):

that is the answer to question one

OpenStudy (misty1212):

ready for part two?

OpenStudy (sissyedgar):

not yet, i have to take away what I had for one then put in whats rite! :P

OpenStudy (misty1212):

kk

OpenStudy (misty1212):

take your time and hurry up

OpenStudy (sissyedgar):

ok! lol, i done

OpenStudy (misty1212):

ok ready?

OpenStudy (sissyedgar):

yes ma'am!

OpenStudy (misty1212):

deposit another quarter

OpenStudy (misty1212):

for the next one we have to divide

OpenStudy (sissyedgar):

quick question

OpenStudy (misty1212):

\[\huge \frac{2.5\times 10^4}{1.578\times 10^{-5}}\]

OpenStudy (misty1212):

ok

OpenStudy (sissyedgar):

did we round the 1st to the nearest 10th?

OpenStudy (misty1212):

we rounded a bit, perhaps we should round some more

OpenStudy (misty1212):

we can't really round to the nearest tenth, i think that is a typo our number had a zero in the tenths place

OpenStudy (sissyedgar):

ooo, ok

OpenStudy (misty1212):

maybe they meant round this \[\huge 1.578\times 10^{-5}\] to this \[\huge 1.6\times 10^{-5}\]

OpenStudy (misty1212):

that would make life easier

OpenStudy (sissyedgar):

mkay

OpenStudy (sissyedgar):

lol

OpenStudy (misty1212):

then the second question we can do \[\frac{2.5\times10^4}{1.6\times 10^{-5}}\]

OpenStudy (misty1212):

\[2.5\div 1.6=1.5625\] and \[\frac{10^4}{10^{-5}}=10^{4-(-5)}=10^9\]

OpenStudy (misty1212):

so we have \[1.5625\times 10^9\] as our answer again maybe you should round or you can leave it like that

OpenStudy (sissyedgar):

lets round... it said round again

OpenStudy (sissyedgar):

1.6 x 10^9

OpenStudy (sissyedgar):

right?

OpenStudy (misty1212):

ok

OpenStudy (misty1212):

that is the answer to number 2

OpenStudy (misty1212):

a billion has 9 zeros

OpenStudy (sissyedgar):

okay... 1 more! YESS! :P

OpenStudy (misty1212):

so 7.046 billion is \(7.046\times 10^9\) that is bigger than \(1.6\times 10^9\)

OpenStudy (misty1212):

oh wait i misread the question sorry

OpenStudy (sissyedgar):

rite

OpenStudy (sissyedgar):

o, ok, lol

OpenStudy (misty1212):

last job it so divide again \[\frac{7.046\times 10^9}{1.6\times 10^9}\]

OpenStudy (misty1212):

about \(4.4\)

OpenStudy (misty1212):

and now we are done!!

OpenStudy (sissyedgar):

o, oky! TY MISTY!!!!!!!!!!!!! I wont loose it this time, i wrote it as u did! :P ~Hugz~

OpenStudy (misty1212):

\[\huge \color\magenta\heartsuit\]